Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Saturday, November 14, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 14, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/14/98 (Saturday)

Segment 1 (Emidio to ANF) Report for 14 November, 1998. Excavation of pit at the Culverts on Peace Valley Road (EMP 25.45) continued. Backfill operations worked from EMP 7.07, Sabodan Road, along Access Road B to EMP 6.0. On the Tejon Bench, the Tie-in at EMP 18.51 was completed and backfill of the pipe was started. Excavation continued at EMP 18.12, where a segment of pipe was lowered in to ditch yesterday. Excavation on Grapevine road continued to ~EMP 15.70. The Bending Machine is set up at EMP 16.10, as pipe segments are lowered in just north of that location. Welding progressed to ~EMP 16.20. Welding and Electrical crews continued work at Grapevine station with some work taking place at Emidio Station. At the California aqueduct bore site, they have progressed approximately 115' of the 400' to be bored. The operation was "back boring" today. At the Tecuya Creek and I-5 bore site, pipe continued to be pushed through casing. Shading and Fiber optic Conduit was placed from ~EMP 17.50 to `~EMP 17.32.

Segment 2 (ANF to Sylmar) Report for 11/14/98: 2C (Templin Hwy to Castaic): Recontouring continued south of Templin Hwy, EMP 50.89 to 50.92, at the site of the testing cutout. The only other work on Segment 2C today was Recontouring in Marple Cyn north to EMP 55.87. 2B (Castaic to Sylmar): Pulling of F/O cable continued on The Old south of Hasley Cyn. Boring continued on the Santa Clara River directional drill at Feedmill Rd. Backfill began at the Calgrove bore site. Between Weldon Cyn and Sierra Hwy, sawcutting in the roadway, trenching, and lowering-in of pipe took place. The bending operation continued just south of Sierra Hwy. At Sunshine Cyn, slurring, shading, and paving occurred.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 14 November 1998: 3A: Backfill crews continued their work on the MTA right of way between Tyburn and Fletcher, filling in sink holes and assist pits. Fiber optic crews blew rope into the conduit between Chevy Chase and Highland Drive. The segment 3 yard in Glendale was moved to the San Fernando Yard.

Segment 4B(El Segundo to Lynwood) 11-14-98: The crew at El Segundo Station ESEMP 11.00 continued to work on the 90 degree tie-in. The crew at Western ESEMP 5.56 was coating and backfilling. The crew at Lynwood station was sending the cleaning pigs to the Corn field yard. The crew inside Lynwood station were working on the electrical wires and the back wall. The crew at College St. was repairing the fiberoptic conduit.

Segment 5 (Lynwood to Long Beach) Report for 14 November 1998: The ream was complete at the Highway 91/Compton Creek directional drill. The pot-hole crew for the Wilmington tie-in encountered contaminated soil. Soils were excavated and stored separately and covered until disposal is determined. The bore crew continued with cased bore into refinery.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page